OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using elimination: -7x + 5y = -12 8y − 3x + 3 = 0 A. Write the second equation in terms of x. B. Replace the right side of the first equation with the second equation. C. Multiply the second equation by -7. D. Multiply the first equation by -5.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@tvcong : whys that ?

OpenStudy (anonymous):

You should try it to see if it is the first step.

OpenStudy (anonymous):

i dont understand what writing the second equation in te

OpenStudy (anonymous):

*terms of x means

OpenStudy (anonymous):

I think it means solving for y, and y is expressed in terms of x.

OpenStudy (richyw):

I would multiply the second equation by -7 as the first step. From there you could multiply the first equation by three and subtract them to eliminate x from one giving you y which you can use to get x.

OpenStudy (anonymous):

A. Write the second equation in terms of x. it mean write rewrite equtionn like this -7x + 5y = -12 8y − 3x +=-3

OpenStudy (richyw):

put into an augmented matrix \[\left[\begin{matrix}-7 & 5 & -12 \\ -3 & 8 & -3\end{matrix}\right]\] multiply second row by -7 \[\left[\begin{matrix}-7 & 5 & -12 \\ 21 & -56 & 21\end{matrix}\right]\] multiply first row by -3 \[\left[\begin{matrix}21 & -15 & 36 \\ 21 & -56 & 21\end{matrix}\right]\] subtract second from first \[\left[\begin{matrix}21 & -15 & 36 \\ 0 & 41 & 15\end{matrix}\right]\] so 41y=15 ==> y=15/41 and 21x-15y=36 21x-15*15/41=36 21x=1704/41 ==> x=81/41 so the solutions are x=81/41 and y=15/41. As you can see, the first step I took was multiplying the second equation by negative 7, and I was able to solve the equation without ever dividing it by 7 (you can start with ANY of these and then "undo" them later and arrive at the solution), therefore the answer "C" is satisfied

OpenStudy (richyw):

the method I used is called the "method of elimination" so C is definitely the answer